# AceCoder: Acing Coder RL via Automated Test-Case Synthesis

### v3

@@ 1 Introduction
In recent years, code generation models have advanced significantly with compute scaling (Kaplan et al., 2020) and training data quality improvement (Huang et al., 2024; Lozhkov et al., 2024; Guo et al., 2024b). The state-of-the-art coder models, including Code-Llama (Rozière et al., 2023), Qwen2.5-Coder (Hui et al., 2024a), DeepSeek-Coder (Guo et al., 2024a) and so on, have shown unprecedented performance across a wide range of coding tasks like program synthesis (Chen et al., 2021), program repair (Zheng et al., 2024a), optimization (Shypula et al., 2023), test generation (Steenhoek et al., 2023), SQL (Yu et al., 2018), issue fix (Jimenez et al., 2024). These models are all pre-trained and further supervised fine-tuned (SFT) on large-scale coding data from web resources like Common Crawl or Github.
@@
(2) Scarcity of large-scale coding datasets with reliable test cases. Most existing coding datasets like APPS (Hendrycks et al., 2021; Chen et al., 2021) heavily rely on costly human expert annotations for test cases, which limits their scalability for training purposes. The largest data is TACO (Li et al., 2023) with 25K examples, which are crawled from the popular coding competition websites, which were already heavily exploited during the pre-training phase.
@@ 2 Methodology
In this section, we will introduce the overall methodology of AceCoder. We begin with formulations of the problems we are investigating, including reward model training and reinforcement learning for LLMs. We then elaborate on how we synthesize the test cases and construct the AceCode-87K. Finally, we explain how we perform the reinforcement learning using our AceCode-RM trained on the AceCode-87K.
@@ 2.1 Problem Formulation
After we get the trained reward model $R_{\phi}$, one way to quickly test the performance of the reward model is Best-of-N sampling, which is usually used as a test-time scaling approach. We will simply select the best response according to the predicted value of $R_{\phi}$. That is $\mathbf{y}^{*}=\arg\max_{\mathbf{y}^{i}\in{\mathbf{y}^{1},...,\mathbf{y}^{N}}}R_{\phi}(\mathbf{x},\mathbf{y}^{i})$.
@@
We can finally conduct reinforcement learning for the original policy model $\pi_{\theta}$ after we get a well-trained reward model $R_{\phi}$. Proximal Policy Optimization (PPO) is an actor-critic RL algorithm that is widely used for LLM’s RL process. Let $\pi_{\theta_{old}}$ be the reference model and $\pi_{\theta}$ be the current policy model that we are updating frequently during the RL training. We denote $r_{t}(\theta)$ as the probability ratio of the current policy model over the old policy model on the $t$-th generated token:
@@
where $\mathbf{y}\sim\pi_{\theta_{old}}(\cdot|x)$, and $A_{t}$ is the advantage computed through the Generalized Advantage Estimation (GAE) (Schulman et al., 2015) via the rewards generated by $R_{\phi}$ and the learned value function $V_{\psi}$. The PPO training objective will force the policy model $\pi$ to increase the probability of generating tokens with higher $A_{t}$ and decrease the probability ratio of generating tokens with lower $A_{t}$ until the clipped bounds $1+\epsilon$ and $1-\epsilon$ are reached respectively.
However, PPO usually requires training an additional value model $V_{\psi}$ and thus makes the training inefficient. Recently, there are some other works like Reinforecement++ (Hu, 2025) that eliminate the need for value model but instead compute advantage only using the rewards generated by $R_{\phi}$ and the KL-divergence of the tokens after the $t$-th tokens. This makes the RL process more efficient and has also proved to be more stable.
3 AceCode-87K
To be able to train a reward model specifically designed for code generation, the first thing is to synthesize reliable test cases for each coding problem and use them as training signals. In this section, we explain the whole procedure of constructing AceCode-87K step by step. We show the overall statistics in Table 1.
@@ 3 AceCode-87K
We start from existing coding datasets with provided question $\mathbf{x}$ and corresponding program $\mathbf{y}$. Specifically, we combine Magicoder-Evol-Instruct111ise-uiuc/Magicoder-Evol-Instruct-110K, Magicoder-OSS-Instruct-75K222ise-uiuc/Magicoder-OSS-Instruct-75K, and StackPyFunction333bigcode/stack-dedup-python-fns as our seed dataset. We only keep the questions written in Python that contain either a function or a class, resulting in a total of 124K entries. We find that these datasets contain highly noisy questions that could not be easily evaluated using test cases. Therefore, we feed every question-solution pair ($\mathbf{x}$, $\mathbf{y}$) into a GPT-4o-mini (Hurst et al., 2024) to propose a refined LeetCode-style question $\mathbf{x_{r}}$ with highly structured instructions. Meanwhile, we also prompt it to ‘imagine’ around 20 test cases $(t_{1},...,t_{m})$ for each refined coding question $\mathbf{x_{r}}$ based on its understanding of the expected behavior of the desired program. See prompt template used in subsection A.1. Please note that we do not use the program solution $\mathbf{y}$ from the existing datasets at all in our final curated AceCode-87K. These datasets are purely used as seeds to help LLM formulate well-structured coding problems.
@@
We propose to use the Bradley-Terry model to train the reward model as defined in Equation 1. Therefore, we need to construct (question, [positive program, negative program]) data from AceCode-87K. Specifically, we sample programs $(\mathbf{y}^{1},...,\mathbf{y}^{n})$ from existing models (e.g. Llama-3.1 (Grattafiori et al., 2024)) w.r.t $\mathbf{x}_{r}$ and utilize the test-case pass rate to distinguish positive and negative programs. Since the pass rate $s_{i}$ for the sampled program $\mathbf{y}^{i}$ can be any number between $[0,1]$, a minor difference in pass rate may not represent that one program is more accurate than another. Therefore, instead of using $\mathbbm{1}[s_{i}>s_{j}]$ to select the preference pairs, we have thus modified the selection rules to be:
@@
This is to ensure the preferred program has at least a $0.8$ pass rate to make sure it represents a more correct program. Also, we find many sampled programs with 0 pass rates can be caused by some small syntax errors or some Python packaging missing errors during evaluation, we chose to not include them as the preference pair to make sure our constructed datasets represent only the preference-based on the valid pass rate. We also ensure the sampled programs all come from the backbone of $R_{\phi}$ so the reward model is trained in an on-policy way. After that, we train our reward model $R_{\phi}$ by fully fine-tuning an instruct coding model. Specifically, We extract the last token’s final hidden representations and pass it through a linear model head that generates a single scalar output, which is optimized via the loss function defined in Equation 1.
@@ 4.1 Reward Model Training Setup
Table 3: AceCoder’s Performance after RL tuning using Reinforcement++ algorithm. We start with 3 different initial policy models and 2 kind of reward types, where $RM$ means using our trained AceCode-RM and $Rule$ means using the binary pass rate. Results show consistent improvement across various benchmarks.
@@
Our reward model is trained using LlamaFactory (Zheng et al., 2024b). We apply full fine-tuning with DeepSpeed stage 3. We train for 1 epoch using a cosine learning rate schedule, starting at 1e-5 with a warmup ratio of 0.1 to gradually increase the learning rate in the initial training phase. Training batch size is set to 128. We enable bf16 precision to reduce memory overhead without compromising model fidelity. The training takes 24 hours on 8 x A100 GPUs.
@@ 4.2 Reinforcement Learning Setup
We perform RL training from three policy models: Qwen2.5-7B-Instruct  [^5] Qwen/Qwen2.5-7B-Instruct and Qwen2.5-Coder-7B-Base  [^6] Qwen/Qwen2.5-Coder-7B and Qwen2.5-Coder-7B-Instruct. Two types of reward can be used, i.e. the trained reward model AceCode-RM-7B and the rule-based reward, i.e. pass rate over the test cases in AceCode-87K. During training, we set the pass rate to be a binary reward, which is 1.0 when all test cases passed, otherwise 0. This is similar to the verfiable reward used in Tulu3 (Lambert et al., 2024a) and DeepSeek-R1 (Guo et al., 2025). Similar to DeepSeek-R1 (Guo et al., 2025), we also experiment with RL from the base model because SFT may cause the search space of the model to be stuck in the local minimum. Since coding is also a highly verifiable task like math, we include the Qwen2.5-Coder-7B-Base in our experiments.
We have trained different policy model backbones with different rewards, resulting in 6 RL models in total. All the RL-tuning are based on OpenRLHF (Hu et al., 2024). We adopt the Reinforcement++ (Hu, 2025) algorithm instead of PPO to improve the training efficiency without training the value model. It’s also proved to be more stable than PPO and GRPO. We train our model on a subsampled hard version of AceCode-87K, where we keep the 25% of the questions with lower average pass rates and higher variance. This is to ensure the question is hard and that the sampled programs are diverse enough. For the training hyperparameters, we set the rollout batch size to 256, and 8 programs are sampled from per question. The training batch size is 128 with a learning rate of 5e-7. All the models are trained for 1 episode and finished in 6 hours on 8 x H100 GPUs.
@@ 4.3 Evaluation Setup
We evaluate our method on three established code-focused benchmarks: EvalPlus (Liu et al., 2023, 2024b), Big Code Bench (Zhuo et al., 2024) and Live Code Bench (Jain et al., 2024). These benchmarks collectively cover a diverse array of coding tasks, enabling us to assess both the correctness and quality of generated code. For Best-of-N sampling experiments, we adopt top-p sampling with a temperature of 1.0 to generate multiple candidate solutions per question. We select the response with the highest reward for evaluation. For RL experiments, we use the benchmark’s default setting, which is greedy sampling most of the time.
@@ 4.4 Main Results
Here we show the experimental results of reward model and RL-trained model.
@@
We perform RL training over 3 different initial policy models in Table 3 with model-based and rule-based rewards. When starting from Qwen2.5-Instruct-7B, we can see the RL tuning can consistently improve the performance, especially on HumanEval and MBPP. Even for the Plus version with more and harder test cases, the RL-tuned model also has more than $3$ points of improvement.
When starting from the Qwen2.5-Coder-Instruct-7B itself, we can still observe improvements, especially when using the rule-based reward. For example, we get more than $3.4$ improvement on BigCodeBench-Full-Hard. Using the reward model for RL can also bring $3.2$ improvement on MBPP. This highlights the charm of self-improvement given the reward model backbone is the same with the initial policy model. We compare our method with other RL-based models like RLEF (Chen et al., 2024), PPOCoder (Shojaee et al., 2023a), StepCoder (Dou et al., 2024b), DSTC (Liu et al., 2024c), etc. We show that our 7B model is able to beat these competitors across the evaluation benchmarks.
Another experiment we conduct is to perform RL training directly from base model Qwen2.5-Coder-7B-base. We show significant improvement, especially through test-case pass rewards on HumanEval, MBPP, and BigCodeBench-I. These results are achieved by only training for 80 steps. We believe further scaling up the training will lead to much larger gains.
@@
We compare our AceCode-RM with 3 top-ranked RM on the RewardBench, including InternLM2-RM-8B (Cai et al., 2024), Skywork-Llama-3.1-8B, and Skywork-Gemma-27B (Liu et al., 2024a), where results are reported in Table 4. We can see that these general-purpose RM can hardly improve and sometimes decrease the performance through Best-of-N sampling compared to greedy sampling, showcasing the incapability in identifying the correct generated programs. On the other hand, our AceCode-RM surpasses all other publicly released reward models in our evaluation and consistently gets positive gains. These findings further underscore our assumption that previous RM training lacks of reliable signals for codes and prove that our RMs can generate reliable and state-of-the-art reward signals in code generation tasks.

### v4

@@ 1 Introduction
In recent years, code generation models have advanced significantly with compute scaling (Kaplan et al., 2020) and training data quality improvement (Huang et al., 2024; Lozhkov et al., 2024; Guo et al., 2024b). The state-of-the-art coder models, including Code-Llama (Rozière et al., 2023), Qwen2.5-Coder (Hui et al., 2024a), DeepSeek-Coder (Guo et al., 2024a) and so on, have shown unprecedented performance across a wide range of coding tasks like program synthesis (Chen et al., 2021), program repair (Zheng et al., 2024a), optimization (Shypula et al., 2023), test generation (Steenhoek et al., 2023), SQL (Yu et al., 2018), and issue fix (Jimenez et al., 2024). These models are all pre-trained and further supervised fine-tuned (SFT) on large-scale coding data from web resources like Common Crawl or Github.
@@
(2) Scarcity of large-scale coding datasets with reliable test cases. Most existing coding datasets like APPS (Hendrycks et al., 2021; Chen et al., 2021) heavily rely on costly human expert annotations for test cases, which limits their scalability for training purposes. The largest dataset is TACO (Li et al., 2023), containing 25K examples collected from popular coding competition websites, sources that have already been extensively utilized during the pre-training phase.
@@
Additionally, both AceCode-RM-7B and AceCode-RM-32B models demonstrate strong performance on the RM Bench (Liu et al., 2024c), outperforming existing state-of-the-art reward models across various categories. Specifically, AceCode-RM-32B achieves the highest average score of 76.1, leading in key categories such as Coding, Chat, Normal, and Hard tasks. Even the smaller AceCode-RM-7B model shows notable competitiveness, excelling particularly in Coding and Hard categories.
@@ 2 Methodology
In this section, we will introduce the overall methodology of AceCoder. We begin with formulations of the problems we are investigating, including reward model training and reinforcement learning for LLMs. Finally, we elaborate on how we synthesize the test cases and construct the AceCode-87K.
@@ 2.1 Problem Formulation
After we get the trained reward model $R_{\phi}$, one way to quickly test the performance of the reward model is Best-of-N sampling, which is usually used as a test-time scaling approach. We will simply select the highest-scored response according to the predicted value of $R_{\phi}$. That is $\mathbf{y}^{*}=\arg\max_{\mathbf{y}^{i}\in{\mathbf{y}^{1},...,\mathbf{y}^{N}}}R_{\phi}(\mathbf{x},\mathbf{y}^{i})$.
@@
We can finally conduct reinforcement learning for the original policy model $\pi_{\theta}$ after we get a well-trained reward model $R_{\phi}$. Proximal Policy Optimization (PPO) is an actor-critic RL algorithm that is widely used for LLM’s RL process. Let $\pi_{\theta_{old}}$ be the reference model and $\pi_{\theta}$ be the current policy model that is iteratively updated during the RL training. We denote $r_{t}(\theta)$ as the probability ratio of the current policy model over the old policy model on the $t$-th generated token:
@@
where $\mathbf{y}\sim\pi_{\theta_{old}}(\cdot|x)$, and $A_{t}$ is the advantage computed through the Generalized Advantage Estimation (GAE) (Schulman et al., 2015) via the rewards generated by $R_{\phi}$ and the learned value function $V_{\psi}$. The PPO training objective will force the policy model $\pi$ to increase the probability of generating tokens with higher $A_{t}$ and decrease the probability ratio of generating tokens with lower $A_{t}$ until the clipped bounds $1+\epsilon$ and $1-\epsilon$ are reached, respectively.
However, PPO usually requires training an additional value model $V_{\psi}$ and thus makes the training inefficient. Recently, there have been some other works like Reinforecement++ (Hu, 2025) that eliminate the need for a value model but instead compute advantage only using the rewards generated by $R_{\phi}$ and the KL-divergence of the tokens after the $t$-th token. This makes the RL process more efficient and has also proved to be more stable.
2.2 AceCode-87K
To be able to train a reward model specifically designed for code generation, the first step is to synthesize reliable test cases for each coding problem and use them as training signals. In this section, we explain the whole procedure of constructing AceCode-87K step by step. We show the overall statistics in Table 1.
@@ 2.2 AceCode-87K
We start from existing coding datasets with provided question $\mathbf{x}$ and corresponding program $\mathbf{y}$. Specifically, we combine Magicoder-Evol-Instruct111ise-uiuc/Magicoder-Evol-Instruct-110K, Magicoder-OSS-Instruct-75K222ise-uiuc/Magicoder-OSS-Instruct-75K, and StackPyFunction333bigcode/stack-dedup-python-fns as our seed dataset. We only keep the questions written in Python that contain either a function or a class, resulting in a total of 124K entries. We find that these datasets contain highly noisy questions that cannot be easily evaluated using test cases. Therefore, we feed every question-solution pair ($\mathbf{x}$, $\mathbf{y}$) into a GPT-4o-mini (Hurst et al., 2024) to propose a refined LeetCode-style question $\mathbf{x_{r}}$ with highly structured instructions. Meanwhile, we also prompt it to ‘imagine’ around 20 test cases $(t_{1},...,t_{m})$ for each refined coding question $\mathbf{x_{r}}$ based on its understanding of the expected behavior of the desired program. See prompt template used in subsection A.3. Please note that we do not use the program solution $\mathbf{y}$ from the existing datasets at all in our final curated AceCode-87K. These datasets are purely used as seeds to help LLM formulate well-structured coding problems.
@@
Human Case Study
Finally, as a last verification step, we conducted a human case study by randomly selecting 40 questions, each with 5 randomly sampled test cases, resulting in 200 manually annotated test cases. Encouragingly, only 3 out of these 200 test cases were found invalid upon review, indicating that our filtering methodology effectively mitigates hallucinations and preserves test-case quality. Moreover, Table 2 presents various pass-rate statistics for AceCode-87K, illustrating the overall difficulty and performance trends.
Pass @ 1 34.90%
Pass @ 4 38.50%
Pass @ 8 39.20%
Pass @ 16 39.60%
Avg Test Case Pass % 70.30%
% Question Where All 16 Inferences Pass All Test Cases 22.80%
Table 2: Dataset pass-rate statistics of AceCode-87K after test-case filtering using Qwen2.5-Coder-7B-Instruct (16 responses per query with temperature=1.0)
@@
We propose to use the Bradley-Terry loss to train the reward model as defined in Equation 1. Therefore, we need to construct (question, [positive program, negative program]) data from AceCode-87K. Specifically, we sample programs $(\mathbf{y}^{1},...,\mathbf{y}^{n})$ from existing models (e.g. Llama-3.1 (Grattafiori et al., 2024)) w.r.t $\mathbf{x}_{r}$ and utilize the test-case pass rate to distinguish positive and negative programs. [^4] Appendix A.2 contains a detailed model breakdown for how different models are used throughout the training and evaluation process of AceCode-RM for more clarifications. Since the pass rate $s_{i}$ for the sampled program $\mathbf{y}^{i}$ can be any number between $[0,1]$, a minor difference in pass rate may not represent that one program is more accurate than another. Therefore, instead of using $\mathbbm{1}[s_{i}>s_{j}]$ to select the preference pairs, we have thus modified the selection rules to be:
@@
This is to ensure the preferred program has at least a $0.8$ pass rate to make sure it represents a more correct program. Also, we find that many sampled programs with 0 pass rates can be caused by some small syntax errors or some Python packaging errors during evaluation. We chose not to include them as the preference pair to make sure our constructed datasets represent only the preference based on the valid pass rate. We also ensure that the sampled programs all come from the backbone of $R_{\phi}$, so the reward model is trained in an on-policy way. After that, we train our reward model $R_{\phi}$ by fully fine-tuning an instruction coding model. Specifically, we extract the last token’s final hidden representation and pass it through a linear model head that generates a single scalar output, which is optimized via the loss function defined in Equation 1.
@@ 3.1 Reward Model Training Setup
Table 4: AceCoder’s Performance after RL tuning using Reinforcement++ algorithm. We start with 3 different initial policy models and 2 kinds of reward types, where $RM$ means using our trained AceCode-RM and $Rule$ means using the binary pass rate. Results show consistent improvement across various benchmarks.
Our reward model is trained using LlamaFactory (Zheng et al., 2024b). We apply full fine-tuning with DeepSpeed stage 3. We train for 1 epoch using a cosine learning rate schedule, starting at 1e-5 with a warmup ratio of 0.1 to gradually increase the learning rate in the initial training phase. The training batch size is set to 128. We enable BF16 precision to reduce memory overhead without compromising model fidelity. The training takes 24 hours on 8 x A100 GPUs.
@@ 3.2 Reinforcement Learning Setup
We perform RL training from three policy models: Qwen2.5-7B-Instruct  [^6] Qwen/Qwen2.5-7B-Instruct, Qwen2.5-Coder-7B-Base  [^7] Qwen/Qwen2.5-Coder-7B, and Qwen2.5-Coder-7B-Instruct. Two types of reward can be used, i.e. the trained reward model AceCode-RM-7B and the rule-based reward, i.e. pass rate over the test cases in AceCode-87K. During training, we set the pass rate to be a binary reward, which is 1.0 when all test cases pass, otherwise 0. This is similar to the verifiable reward used in Tulu3 (Lambert et al., 2024a) and DeepSeek-R1 (Guo et al., 2025). Similar to DeepSeek-R1 (Guo et al., 2025), we also experiment with RL from the base model because SFT may cause the search space of the model to be stuck in a local minimum. Since coding is also a highly verifiable task like math, we include the Qwen2.5-Coder-7B-Base in our experiments.
We have trained different policy model backbones with different rewards, resulting in 6 RL models in total. All the RL-tuning is based on OpenRLHF (Hu et al., 2024). We adopt the Reinforcement++ (Hu, 2025) algorithm instead of PPO to improve the training efficiency without training the value model. It’s also proved to be more stable than PPO and GRPO. We train our model on a subsampled hard version of AceCode-87K, where we keep the 25% of the questions with lower average pass rates and higher variance. This is to ensure the question is hard and that the sampled programs are diverse enough. For the training hyperparameters, we set the rollout batch size to 256, and 8 programs are sampled per question. The training batch size is 128 with a learning rate of 5e-7. All the models are trained for 1 episode and finished in 6 hours on 8 x H100 GPUs.
@@ 3.3 Evaluation Setup
We evaluate our method on four established code-focused benchmarks: HumanEval(+) (Chen et al., 2021; Liu et al., 2023), MBPP(+) (Austin et al., 2021; Liu et al., 2023), BigCodeBench (Zhuo et al., 2024) and LiveCodeBench (V4) (Jain et al., 2024). These benchmarks collectively cover a diverse array of coding tasks, enabling us to assess both the correctness and quality of generated code. For Best-of-N sampling, we adopt top-p sampling with a temperature of 1.0 to generate multiple (16/32/64) candidate solutions per question and then select the response with the highest reward for evaluation. For RL experiments, we use each benchmark’s default setting, which is greedy sampling most of the time.
@@ 3.4 Main Results
Here we show the experimental results of the reward models and RL-trained models.
@@
We perform RL training over 3 different initial policy models in Table 4 with model-based and rule-based rewards. When starting from Qwen2.5-Instruct-7B, we can see that RL tuning can consistently improve performance, especially for HumanEval and MBPP. Even for the Plus version with more and harder test cases, the RL-tuned model also has more than $3$ points of improvement.
When starting from the Qwen2.5-Coder-Instruct-7B itself, we can still observe improvements, especially when using the rule-based reward. For example, we get more than $3.4$ improvement on BigCodeBench-Full-Hard. Using the reward model for RL can also bring a $3.2$ improvement on MBPP. This highlights the charm of self-improvement given the reward model backbone is the same as the initial policy model. We compare our method with other RL-based models like RLEF (Chen et al., 2024), PPOCoder (Shojaee et al., 2023a), StepCoder (Dou et al., 2024b), DSTC (Liu et al., 2024d), etc. We show that our 7B model is able to beat these competitors across the evaluation benchmarks.
Another experiment we conduct is to perform RL training directly from the base model Qwen2.5-Coder-7B-base. We show significant improvement, especially through test-case pass rewards on HumanEval, MBPP, and BigCodeBench-I. These results are achieved by only training for 80 steps. We believe further scaling up the training will lead to much larger gains.
@@
We compare our AceCode-RM with 3 top-ranked RM on the RewardBench, including InternLM2-RM-8B (Cai et al., 2024), Skywork-Llama-3.1-8B, and Skywork-Gemma-27B (Liu et al., 2024a), where results are reported in Table 5. We can see that these general-purpose RM can hardly improve and sometimes decrease the performance through Best-of-N sampling compared to greedy sampling, showcasing the incapability in identifying the correct generated programs. On the other hand, our AceCode-RM surpasses all other publicly released reward models in our evaluation and consistently gets positive gains. These findings further underscore our assumption that previous RM training lacks reliable signals for codes and prove that our RMs can generate reliable and state-of-the-art reward signals in code generation tasks.
Model Code Chat Math Safety Easy Normal Hard Avg
Skywork/Skywork-Reward-Llama-3.1-8B 54.5 69.5 60.6 95.7 89 74.7 46.6 70.1
NVIDIA/Nemotron-340B-Reward 59.4 71.2 59.8 87.5 81 71.4 56.1 69.5
internlm/internlm2-20b-reward 56.7 63.1 66.8 86.5 82.6 71.6 50.7 68.3
internlm/internlm2-7b-reward 49.7 61.7 71.4 85.5 85.4 70.7 45.1 67.1
Skywork-Reward-Llama-3.1-8B-v0.21 53.4 69.2 62.1 96 88.5 74 47.9 70.1
Skywork-Reward-Gemma-2-27B-v0.21 45.8 49.4 50.7 48.2 50.3 48.2 47 48.5
AceCodeRM-7B 66.9 66.7 65.3 89.9 79.9 74.4 62.2 72.2
AceCodeRM-32B 72.1 73.7 70.5 88 84.5 78.3 65.5 76.1
$\Delta$ (AceCodeRM-7B - max(other RM)) 7.5 -4.5 -6.1 -6.1 -9.1 -0.3 6.1 2.1
$\Delta$ (AceCodeRM-32B - max(other RM)) 12.7 2.5 -0.9 -8 -4.5 3.6 9.4 6
  • 1
    There is no official result for this model; however, the authors made best efforts to extend the original RM Bench code base to adapt to this new model.
Table 6: AceCode-RM’s Performance on RM Bench against various other reward models. We can see that AceCode-RM-32B model performed best in Coding, Chat, Normal, Hard, and Average scores against all other reward models.
Moreover, we have also evaluated our AceCode-RM against other reward models on RM-Bench in Table 6, a benchmark designed to assess the reward model’s capabilities across code synthesis, mathematical reasoning, and other tasks (Liu et al., 2024c). In this head-to-head comparison, AceCode-RM emerges as the clear state-of-the-art in coding, hard tasks, and overall average. Remarkably, our 7 billion-parameter variant, AceCode-RM-7B, outperforms NVIDIA-Nemotron-340B-Reward(Nvidia et al., 2024) by 7.50 points on the coding benchmark, proving that a more compact model can deliver superior reward estimates for code generation. Beyond code, AceCode-RM also generalizes well: it beats all other reward models on the average score, underscoring its robust reasoning and dialogue capabilities.